# Traversable wormhole sourced by a generalized string fluid in the 4D Einstein-Gauss-Bonnet gravity

### v1

@@ Abstract
We investigate traversable wormhole solutions in four-dimensional Einstein-Gauss-Bonnet (EGB) gravity coupled with a generalized string fluid background. Our analysis reveals that zero-tidal solutions meet all geometric criteria for traversability while avoiding singularities, with the Gauss-Bonnet coupling $\alpha$ modulating the throat geometry and asymptotic behavior. Varying the parameters $\alpha$ and $\varepsilon$ significantly affects curvature both near the throat and at large distances, maintaining finite values throughout spacetime. We demonstrate that the null energy condition (NEQ) can be satisfied for high values of $\alpha$ ($\gtrsim 1$) and low string fluid density $\varepsilon$ ($\lesssim 0.1$), greatly reducing the exotic matter requirement compared to general relativistic wormholes. The string fluid’s behavior approaches that of a cloud of strings for large $\alpha$ or $\varepsilon$. Additionally, the volume integral quantifier and complexity factor analyses show that increasing $\alpha$ reduces gravitational complexity and exotic matter content near the throat, aligning with NEQ preservation. Our results establish that the combined influence of Gauss-Bonnet gravity and string-like matter can support traversable wormholes while minimizing unphysical energy conditions, offering new insights for constructing viable wormhole geometries in modified gravity.
@@ I Introduction
Related to the above, for a long time it was assumed that the Gauss-Bonnet Lagrangian in $d\leq 4$ dimensions does not contribute to the equations of motion, since it corresponds to a total derivative. Thus, in $4D$, this Lagrangian was usually treated as a topological invariant. Consequently, equations of motion in $4D$ for Einstein-Gauss-Bonnet (EGB) gravity only made physical sense in the presence of scalar fields coupled to the aforementioned Lagrangian. Recently, in 2020, in reference [23], it was shown that by re-scaling the coupling constant associated with the Gauss-Bonnet term as ${\alpha_{2}}\to\alpha/(d-4)$, the Lanczos tensor remains finite and non-zero. As claimed in that reference, this prescription is analyzed from the perspective of classical field theory. In this regard, it is worth mentioning that, as references [24, 25] point out, the existence of solutions in the $4D$ EGB limit does not necessarily imply the existence of a well-defined theory in four dimensions. Several objections have been raised against the $4D$ EGB case [26, 27, 28]. However, the conclusion that a four-dimensional Gauss-Bonnet gravity theory does not exist turned out to be premature, as argued in references [24, 25]. It was demonstrated that it is indeed possible to take the limit $D\to 4$ in the action of EGB [29, 30], generalizing a previous procedure used to take the limit $D\to 2$ in general relativity [31]. Wormhole solutions in 4D EGB gravity have also been explored in the literature. See, for example references [32, 33, 34, 35, 36, 37].
@@ III.4 Generalized string fluid described by a Kiselev-like source
This relation reveals how $\alpha$ must scale with $r_{0}$ in order to maintain $\omega_{q}(r_{0})=-1$. Figure 6 (left panel) illustrates this for $\varepsilon=0.1$ and $a=1$. The right panel of Fig. 6 confirms that this parameter choice indeed yields $\omega_{q}(r_{0})=-1$, while asymptotically approaching $\omega_{q}\rightarrow-1/3$ (cosmic string behavior, dashed line). Notably, this asymptotically approaching limit aligns with observational constrains on late-time cosmic acceleration [3, 8].
@@ IV Null Energy Condition and Quantity of Exotic Matter
This expression shows that, for fixed values of $\varepsilon$, $r_{0}$, and $a$, the VIQ evaluated near the wormhole throat can be made arbitrarily small by increasing the coupling constant $\alpha$. This suggests that the total amount of exotic matter required to sustain the wormhole may become negligible in the strong-coupling regime of Einstein-Gauss-Bonnet gravity, as already demonstrated in the previous analysis concerning the radial state parameter.
@@ VI Conclusion
Acknowledgments. This work was supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Fundação Cearense de Apoio ao Desenvolvimento Científico e Tecnológico (FUNCAP). LCNS would like to thank FAPESC for financial support under Grant No. 735/2024. The authors thank colleagues at Universidade Estadual do Ceará for valuable discussions.

### v2

@@ Abstract
We investigate traversable wormhole solutions in four-dimensional Einstein-Gauss-Bonnet (EGB) gravity coupled with a generalized string fluid background. Our analysis reveals that zero-tidal solutions meet all geometric criteria for traversability while avoiding singularities, with the Gauss-Bonnet coupling $\alpha$ modulating the throat geometry and asymptotic behavior. Varying the parameters $\alpha$ and $\varepsilon$ significantly affects curvature both near the throat and at large distances, maintaining finite values throughout spacetime. We demonstrate that the null energy condition (NEC) can be satisfied for high values of $\alpha$ ($\gtrsim 1$) and low string fluid density $\varepsilon$ ($\lesssim 0.1$), greatly reducing the exotic matter requirement compared to general relativistic wormholes. The string fluid’s behavior approaches that of a cloud of strings for large $\alpha$ or $\varepsilon$. Additionally, the volume integral quantifier and complexity factor analyses show that increasing $\alpha$ reduces gravitational complexity and exotic matter content near the throat, aligning with NEC preservation. Our results establish that the combined influence of Gauss-Bonnet gravity and string-like matter can support traversable wormholes while minimizing unphysical energy conditions, offering new insights for constructing viable wormhole geometries in modified gravity.
@@ I Introduction
Related to the above, for a long time it was assumed that the Gauss-Bonnet Lagrangian in $d\leq 4$ dimensions was non-dynamical since it represents a topological invariant (total derivative) and does not contribute to the equations of motion. Consequently, equations of motion in $4D$ for Einstein-Gauss-Bonnet (EGB) gravity only are physically meaningful in the presence of scalar fields coupled to the aforementioned Lagrangian. Recently, in 2020, in reference [23], it was shown that by re-scaling the coupling constant associated with the Gauss-Bonnet term as ${\alpha_{2}}\to\alpha/(d-4)$, the Lanczos tensor remains finite and non-zero. However, it seems that the model faces consistency challenges in topological actions and divergent entropy. Lü & Pang [24] addressed these issues through dimensional compactification, showing that the $D\rightarrow 4$ limit equivalently describes a Horndeski theory with an additional scalar field. This duality suggests that solutions in 4D EGB - such as black holes [23] or wormholes - may carry hidden scalar degrees of freedom if analyzed from the perspective of classical field theory.
In that regard, it is worth mentioning that, as references [25, 26] point out, the existence of solutions in the $4D$ EGB limit does not necessarily imply the existence of a well-defined theory in four dimensions. Several objections have been raised against the $4D$ EGB case [27, 28, 29, 30]. However, the conclusion that a four-dimensional Gauss-Bonnet gravity theory does not exist turned out to be premature, as argued in references [25, 26]. It was demonstrated that it is indeed possible to take the limit $D\to 4$ in the action of EGB [31, 32], generalizing a previous procedure used to take the limit $D\to 2$ in general relativity [33]. Wormhole solutions in 4D EGB gravity have also been explored in the literature. See, for example references [34, 35, 36, 37, 38, 39]. Additionally, black hole configurations in the same regularized 4D EGB framework have been investigated under scalar and particle field perturbations, particularly in the context of the weak cosmic censorship conjecture [40].
@@ III.4 Generalized string fluid described by a Kiselev-like source
This relation reveals how $\alpha$ must scale with $r_{0}$ in order to maintain $\omega_{q}(r_{0})=-1$. Figure 6 (left panel) illustrates this for $\varepsilon=0.1$ and $a=1$. The right panel of Fig. 6 confirms that this parameter choice indeed yields $\omega_{q}(r_{0})=-1$, while asymptotically approaching $\omega_{q}\rightarrow-1/3$ (cosmic string behavior, dashed line). Notably, this asymptotically approaching limit aligns with observational constrains on late-time cosmic acceleration [3, 8, 48].
@@ IV Null Energy Condition and Quantity of Exotic Matter
This expression shows that, for fixed values of $\varepsilon$, $r_{0}$, and $a$, the VIQ evaluated near the wormhole throat can be made arbitrarily small by increasing the coupling constant $\alpha$. In contrast to conventional solutions that require violation of the NEC, or even models with frozen scalar fields [51], our configuration with generalized string fluid can therefore preserve the NEC for certain ranges of $\alpha$ and $\varepsilon$, reducing the need for exotic matter and suggesting that the total amount of exotic matter required to sustain the wormhole may become negligible in the strong-coupling regime of Einstein-Gauss-Bonnet gravity, as already demonstrated in the previous analysis concerning the radial state parameter.
@@ VI Conclusion
Acknowledgments. This work was partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under Grants 308268/2021-6 (CRM) and 315926/2021-0 (MSC). ME is funded by ANID , FONDECYT de Iniciación en Investigación 2023, Folio 11230247. LCNS would like to thank FAPESC for financial support under Grant No. 735/2024. The authors thank colleagues at Universidade Estadual do Ceará for valuable discussions.